OpenStudy (anonymous):

Does this online tutoring thing really work? I am studying for the GRE and have math questions!

OpenStudy (turingtest):

Ask a question and find out :)

OpenStudy (anonymous):

One of the roots of the equation x^2 + kx -6 = 0 is 3, and k is a constant. Quantity A: The value of k QUantity B: -1 Answer choices: QUantity A is greater Quantity B is greater THe two quantitis are equal Relationship cannot be determined from given info

OpenStudy (radar):

I THINK I GOOFED. ONE OF THE ROOTS IS A 3. (x-3)(x+2)= \[x ^{2}-x-6=0\]

OpenStudy (radar):

So k is a -1 And it is said that Quantitu B=-1 Now the relationship may be more obvious to you.

OpenStudy (anonymous):

In my original question it says that one of the roots is 3, so the way I see it the answer would look something like this: (x-3)(x+2)=0 factored out it would be: x^2-(1)x-6=0 so I guess the two quantities are equal...? It just looked wrong to me because the original equation they gave was: x^2 PLUS kx -6=0, so that tripped me up.

OpenStudy (anonymous):

oh yup, we both goofed it the first time we tried. these stupid GRE questions arent that hard, its almost more about reading comprehension than anything else. thanks for your help!

OpenStudy (radar):

I believe you have solved it. Great, I gotta go now (take a cat to the vet)